OpenStudy (daenio):

A drawer contains 13 socks - 4 blue, 3 red, and 6 green. If four socks are drawn randomly (without replacement), the probability that three of them are green is approximately Question 2 options: a) 21.18% b) 19.58%

Parth (parthkohli):

I'll let Mr. Lgba answer on this one =)

OpenStudy (lgbasallote):

lol no..i dont know probabilities

OpenStudy (lgbasallote):

well i used to know them..not anymore

Parth (parthkohli):

Okay, let me answer then :P

Parth (parthkohli):

What is the probability of getting a green sock?

OpenStudy (daenio):

I guess it would be 6/13= 0.46%.

OpenStudy (anonymous):

\( \frac{6}{13} \cdot \frac{5}{12} \cdot \frac{4}{11} \cdot \frac{7}{12}\) is the probability of choosing Green, Green, Green, not Green.

OpenStudy (daenio):

@ParthKohli

Parth (parthkohli):

He's right lol

OpenStudy (daenio):

Where did 7/12 come from?

Parth (parthkohli):

7/12 is the probability of NOT choosing a green one.

Parth (parthkohli):

Since we have to choose it 3 times, not 4.

OpenStudy (kropot72):

\[\frac{\left(\begin{matrix}6 \\ 3\end{matrix}\right)\left(\begin{matrix}7 \\ 1\end{matrix}\right)}{\left(\begin{matrix}13 \\ 4\end{matrix}\right)}\] =0.1958

OpenStudy (daenio):

@ParthKohli I'm a bit confused on why the denominator is 12. (7/12)

OpenStudy (kropot72):

The hypergeometric distribution applies to this problem. See above for solution.

OpenStudy (daenio):

Thank you @kropot72! :)

OpenStudy (kropot72):

You're welcome :)